Whitney's lung capacity was measured as 2.9 L at a body temperature of 37 ∘C and a pressure of 742 mmHg.

How many grams of nitrogen are in Whitney's lungs at STP if air contains 78 % nitrogen?
Answer #1

The moles of air are calculated:

n = P * V / R * T = (742/760) * 2.9 / 0.082 * (37 + 273) = 0.111 mol

The nitrogen mass is calculated:

m N2 = 0.111 mol air * (78 mol N2 / 100 mol air) * (28 g N2 / 1 mol) = 2.42 g

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.